OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). A. y = 3x - 5 B. y = 3x - 2 C. y = 3x + 1 D. y = 3x + 3

OpenStudy (anonymous):

we have given (x1 , y1) = (1 , -2) & slope m = 3 So use slope intercept form y - y1 = m (x - x1) just plug the values & Solve

OpenStudy (sidsiddhartha):

hint: if slope of a line passing through (x,y) is m then equation of that line is y=mx+c c=intercept